OpenStudy (anonymous):

Can anyone please explain how to solve this problem? The sum of a number and its square is 72. Find the number.

OpenStudy (anonymous):

Is this the correct formula to start : n + n^2 = 72

OpenStudy (anonymous):

ya.....let the number be n n+n^2=72 n^2+n-72=0 n^2+8n-9n-72=0 n(n+8)-9(n+8)=0 (n+8)(n-9)=0

OpenStudy (anonymous):

n=9 and -8 value is rejected.....

OpenStudy (anonymous):

Thank you!!!

OpenStudy (phi):

I would tweak the answers. (n+8)(n-9) is not n^2 +n -72 And I don't see any reason to exclude a negative number as a possible answer.